Problems for the Adept

Lesson 4

Simple exercises

• Get a bitmap from somewhere and print it out. Use any bitmap you can find; if you can’t find any, there are at least three available as /mit/postscript/*.bitmap [These are each 32 x 32 bit images].

• Fill in a spiral: Divide the spiral into appropriate divisions and shade them to taste.

Slightly harder

• Look at the items (keys) in a font dictionary and print them out.

• Define a font in which one character in the font is a spiral.

• Write a string in which you fill the characters with smaller version of the character itself (e.g. if you draw an A, it is filled with or built out of smaller A’s).

Still harder …

• Redefine a standard font to have an accented e . (.e.g. é)

• Write text in a spiral. Perhaps have the text diminish and/or fade as it nears the center.

• Write a string that darkens from left to right (yes, the darkness should change within each character).